OpenStudy (anonymous):

Find a44 if a1=-4, d=4! please

jimthompson5910 (jim_thompson5910):

The nth term is an = a1 + (n-1)*d Plug in a1 = -4 and d = 4 to get an = a1 + (n-1)*d an = -4 + (n-1)*4 Now simplify to get the nth term Once you have the nth term, you then plug in n = 44 to find a44

OpenStudy (anonymous):

I got it, I think it is 168:)

jimthompson5910 (jim_thompson5910):

you nailed it